OpenStudy (anonymous):

WILL FAN AN MEDAL Just need my work checked Simplify 9(a+2b)-4(2a-3b) My system : 9a + 18b - 8a +12b = a+ 30b

OpenStudy (anonymous):

if you're just looking for if your work is right it is so far! :)

OpenStudy (anonymous):

is it completely simplified though?

OpenStudy (anonymous):

after you just combine like terms on the left, then subtract =a+30b to the left! :)

OpenStudy (anonymous):

so you would get 9a-8a+= a then 18b+12b= 30b so you have a+30b=a+30b subtract right side to the left side and your answer is gonna be 0!

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

would you mind if i tagged you in my next question ?

OpenStudy (anonymous):

@IsabellaVH